# Viscous-inertial waves on the surface of the Sun: modeling, forward and inverse problems

### v2

@@ 2.1 Reduced-order modeling of inertial waves
with $f:=\rho_{0}\operatorname{curl_{\mathrm{h}}}\textbf{{f}}_{\!\mathrm{h}}$, and $\alpha_{\Omega}$ is a linear function of $\Omega$ as in (2.1) for convenience.
If we are looking for time-harmonic solutions $\Psi(t,\textbf{{r}})=\Re\left(e^{i\omega t}\Psi_{\omega}(\textbf{{r}})\right)$, then the space-dependent part has to satisfy the fourth-order elliptic equation
@@
with $\beta_{\Omega}(\theta):=\Omega(r,\theta)-\Omega_{\rm ref}$.
@@ 2.2 Separated equation as inversion model
Expanding $\Psi_{\omega}(\theta,\phi)=\frac{1}{\sqrt{2\pi}}\sum_{m=-\infty}^{\infty}\widehat{\Psi}_{\omega,m}(\theta)e^{im\phi}$ into a Fourier series in azimuth, the coefficients $\widehat{\Psi}_{\omega,m}$ satisfy the ordinary boundary value problems
$\begin{cases}\gamma\Delta_{m}^{2}\widehat{\Psi}_{\omega,m}(\theta)+i\omega\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)-im\beta_{\Omega}(\theta)\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)+im\alpha_{\Omega}(\theta)\widehat{\Psi}_{\omega,m}(\theta)=\hat{f}_{m}(\theta)\\ (\Gamma_{m}\widehat{\Psi}_{\omega,m})(\theta)=0,\qquad\theta\in P:=\{0,\pi\}\end{cases}$ (9)
@@
$\Gamma_{0}\Psi:=\begin{pmatrix}\Psi^{\prime}|_{P}\\ \Psi^{\prime\prime\prime}|_{P}\end{pmatrix},\quad\Gamma_{\pm 1}\Psi):=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime\prime}|_{P}\end{pmatrix},\quad\Gamma_{m}\Psi:=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime}|_{P}\end{pmatrix},\;|m|\geq 2.$ (10)
We remark that this equation reduces to Eq. (2.9) in [40] if $\gamma=0$ and to Eq. (8) in [12] if $\hat{f}_{m}=0$. However, vanishing Cauchy data are imposed at the poles in [12], which for $|m|\leq 1$ may yield solutions that do not correspond to the unseparated equation (8).
The system (9)-(10) is the underlying model for the parameter inversion problem considered in Section 4. Solving this inverse problem requires measured data, which is the newly discovered inertial waves in this study. The acquisition process will be detailed in the forthcoming section.
2.3 Observation of solar oscillations
The basic input of helioseismology are measurements of line-of-sight velocity based on Doppler shifts of certain spectral lines. These so-called Dopplergrams acquired since 1996 by ground- or space-based telescopes have high-spatial resolution, of at least $1000\times 1000$ pixels, and a temporal cadence of at least $1$ minute. A Fourier transform of the Dopplergram data $v_{\rm dop}(\theta,\phi,t)$ in longitude $\phi$ and time $t$ yields data $v_{\rm dop}^{m,\omega}(\theta)$. Due to the limited data coverage, as only at most half of the Sun is visible at a given time, the transform in $\phi$ can only be integrated on $0,\pi]$ instead of $0,2\pi]$, leading to a leakage of the signal; see, e.g. the review in [4].
Moreover an operator should be applied to convert the simulated $\hat{\Psi}$ into $v_{\rm dop}$ using for example that $v_{\rm dop}=\mathbf{l}\cdot\mathbf{u}$, where $\mathbf{l}$ is the line-of-sight vector. On a formal point of view, this could be written as a (linear) operator that transforms $\Psi$ onto $v_{\rm dop}$. We will not consider this aspect in this article, simply assuming that reprocessed signal $y^{\delta}$ is available and described by
$y^{\delta}(\theta)=\Psi(\theta)+\mathrm{noise}(\theta)\qquad\theta\in I:=(0,\pi)\quad\text{or}\quad\theta\in I^{\prime}:=(\epsilon,\pi-\epsilon),\epsilon\geq 0,$ (11)
here $\delta=\|\mathrm{noise}\|$ denotes the noise level. The case $\epsilon=0$ corresponds to full surface coverage, while $\epsilon>0$ takes into account that higher latitudes cannot easily be observed due to the line-of-sight projection.
It is usually not possible to work with the Doppler signal directly due to the high level of convective noise. However, some inertial modes of large amplitudes have been directly detected [14], enabling us to consider the recovery of viscosity and rotation from the measurements of $y^{\delta}$.
@@ 3 Inertial waves on the manifold $r{\mathbb{S}^{2}}$}
We first show the well-posedness of Eq. (8) associated to the differential operator
@@
with recalling that $\beta_{\Omega}(\theta)=\Omega(\theta)-\Omega_{\mathrm{ref}}$ and $\alpha_{\Omega}(\theta)=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\big{(}\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\big{)}$. We will work in the space $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ of functions in $H^{2}({r{\mathbb{S}^{2}}})$ with zero mean introduced in the Appendix A.
By partial integration with taking into account symmetry of the bi-Laplacian and Laplacian operators (see Appendix, Lemma 19), the corresponding sesquilinear form $\mathbf{A}:H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\times H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\to\mathbb{C}$ defined by $\mathbf{A}[\Psi,\psi]:=\langle\mathbf{B}(\Psi),\psi\rangle_{H_{\diamond}^{-2},H_{\diamond}^{2}}$ is given by
$\begin{aligned} \mathbf{A}[\Psi,\psi]&=\gamma\left\langle\Delta_{\mathrm{h}}\Psi,\Delta_{\mathrm{h}}\psi\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\Psi,\operatorname{\mathbf{grad}_{\mathrm{h}}}\psi\right\rangle+(\mathbf{A}^{\alpha}+\mathbf{A}^{\beta})[\Psi,\psi],\\ &\mathbf{A}^{\alpha}[\Psi,\psi]:=-\left\langle\alpha_{\Omega}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle,\qquad\mathbf{A}^{\beta}[\Psi,\psi]:=\left\langle\beta_{\Omega}\Delta_{\mathrm{h}}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle.\end{aligned}$ (13)
@@
$\quad=\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi_{m},\operatorname{\mathbf{grad}_{\mathit{m}\!}}\psi_{m}\right\rangle+(\mathbf{A}^{\alpha}_{m}+\mathbf{A}^{\beta}_{m})[\Psi_{m},\psi_{m}],$
$\mathbf{A}^{\alpha}_{m}[\Psi_{m},\psi_{m}]:=im\left\langle\alpha_{\Omega}\Psi_{m},\psi_{m}\right\rangle,\qquad\;\mathbf{A}^{\beta}_{m}[\Psi_{m},\psi_{m}]:=-im\left\langle\beta_{\Omega}\Delta_{m}\Psi_{m},\psi_{m}\right\rangle.$
with $\mathbf{A}^{\alpha}_{m},\mathbf{A}^{\beta}_{m}:V_{m}\times V_{m}\to\mathbb{C}$ as restrictions of $\mathbf{A}^{\alpha}$, $\mathbf{A}^{\beta}$ defined in (13). Here, $\Delta_{m}$, $\operatorname{\mathbf{grad}_{\mathit{m}\!}}$ are the decomposed operators as defined in (58), (59) in Appendix A.
@@ 3.1 Boundedness
It is straightforward to show that $\mathbf{A}^{\alpha}$ and $\mathbf{A}^{\alpha}_{m}$ are bounded if $\Omega\in W^{2,\infty}(0,\pi)$. However, we wish to reduce the regularity assumption on the unknown rotation and only assume
@@
Under this reduced regularity assumption, establishing boundedness requires additional technical analysis. For later use in $\mathrm{G\mathring{a}rding}$ inequality, we also provide a bound by a weaker norm. In the following, we will frequently employ continuous embeddings between function spaces $X,Y$ denoted the embedding constants $C_{X\to Y}$, i.e., $\|\cdot\|_{Y}\leq C_{X\to Y}\|\cdot\|_{X}$.
@@ 3.2 Fredholm property and well-posedness
Since the embeddings $H^{1}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ and $H^{3/2}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ are compact, $\mathbf{B}$ is a Fredholm operator of index 0; see, e.g., [33, Sec. 8.2.4].
@@
$\quad\geq\frac{\gamma}{2}\|\Psi\|^{2}_{H_{\diamond}^{2}}+\left(|\omega|\!-\!\frac{4}{\gamma^{3}}(C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}})^{4}\big{(}\|\Omega\!-\!\Omega_{\mathrm{{ref}}}\|^{2}_{H^{1}}\!+\!9\|\Omega\|^{2}_{H^{1}}\big{)}^{2}\right)\|\Psi\|^{2}_{H_{\diamond}^{1}}.$
If $|\omega|$ satisfies (23), then $\mathbf{A}$ is coercive, so $\mathbf{B}$ is boundedly invertible by the Lax-Milgram theorem [11, Section 6.2, Theorem 1]. ∎
@@
The elements of $\operatorname{ker}\mathbf{B}_{\omega}$ for $\omega\in\Lambda$ are called inertial modes and have been studied intensively in [12] and other publications.
@@
Therefore, we are in the framework of the analytic Fredholm theorem (see [33, Thm. 7.92]). Note that $\mathbf{B}_{\omega}$ is boundedly invertible if and only of if $I+\mathbf{C}(\omega)$ is boundedly invertible and that the kernels of both operators coincide. The first alternative of this theory, that is $(I+\mathbf{C}(\omega))^{-1}$ does not exist for any $\omega\in\mathbb{C}$, is excluded by the second part of Theorem 3. Therefore, the second alternative holds true and provides the statements of this corollary. ∎
@@ 3.3 Separated equations
$\left\langle\mathbf{B}_{m}\Psi_{m},\psi_{m}\right\rangle=\mathbf{A}_{m}[\Psi_{m},\psi_{m}],\qquad\Psi_{m},\psi_{m}\in V_{m}.$ (24)
@@
$\quad\leq|m|\left|\Im\big{\langle}\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega}\Psi)\big{\rangle}\right|=|m|\left|\Im\big{\langle}\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega})\Psi\big{\rangle}\right|$
$\quad\leq|m|\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi\|_{L^{6}}\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\beta_{\Omega}\|_{L^{2}}\|\Psi\|_{L^{3}}\leq\frac{|m|}{r}C_{H^{2}_{\diamond}\to L^{3}_{\diamond}}C_{H^{1}_{\diamond}\to L^{6}_{\diamond}}\|\Omega^{\prime}\|_{L^{2}({r{\mathbb{S}^{2}}})}\|\Psi\|_{H^{2}_{\diamond}}^{2}.$

### v3

@@ 2.1 Reduced-order modeling of inertial waves
with $f:=\rho_{0}\operatorname{curl_{\mathrm{h}}}\textbf{{f}}_{\!\mathrm{h}}$, and $\alpha_{\Omega}$ is a linear function of $\Omega$ as in (2.1) for convenience. Lastly, if we are looking for time-harmonic solutions $\Psi(t,\textbf{{r}})=\Re\left(e^{i\omega t}\Psi_{\omega}(\textbf{{r}})\right)$, then the space-dependent part has to satisfy the fourth-order elliptic equation
@@
with $\beta_{\Omega}(\theta):=\Omega(r,\theta)-\Omega_{\rm ref}$. The resulting equation (11) serves as the underlying model of the forward problem studied in Section 3
@@ 2.2 Separated equation as inversion model
To prepare for the inversion analysis, we now derive the separated version of (11) that will serve as our model for parameter identification. Expanding $\Psi_{\omega}(\theta,\phi)=\frac{1}{\sqrt{2\pi}}\sum_{m=-\infty}^{\infty}\widehat{\Psi}_{\omega,m}(\theta)e^{im\phi}$ into a Fourier series in azimuth, the coefficients $\widehat{\Psi}_{\omega,m}$ satisfy the ordinary boundary value problems
$\begin{cases}\gamma\Delta_{m}^{2}\widehat{\Psi}_{\omega,m}(\theta)+i\omega\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)-im\beta_{\Omega}(\theta)\Delta_{m}\widehat{\Psi}_{\omega,m}(\theta)+im\alpha_{\Omega}(\theta)\widehat{\Psi}_{\omega,m}(\theta)=\hat{f}_{m}(\theta)\\ (\Gamma_{m}\widehat{\Psi}_{\omega,m})(\theta)=0,\hskip 18.49988pt\theta\in P:=\{0,\pi\}\end{cases}$ (12)
@@
$\Gamma_{0}\Psi:=\begin{pmatrix}\Psi^{\prime}|_{P}\\ \Psi^{\prime\prime\prime}|_{P}\end{pmatrix},\qquad\Gamma_{\pm 1}\Psi:=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime\prime}|_{P}\end{pmatrix},\qquad\Gamma_{m}\Psi:=\begin{pmatrix}\Psi|_{P}\\ \Psi^{\prime}|_{P}\end{pmatrix},\;|m|\geq 2.$ (13)
We remark that this equation reduces to Eq. (2.9) in [Watson81] if $\gamma=0$ and to Eq. (8) in [Damien22] if $\hat{f}_{m}=0$. However, vanishing Cauchy data are imposed at the poles in [Damien22], which for $|m|\leq 1$ may yield solutions that do not correspond to the unseparated equation (11).
The system (12)–(13) provides the forward model for the parameter inversion problem addressed in Section 4. Solving this inverse problem requires inertial wave observations. The data acquisition process is described in the following section.
2.3 Observation of solar inertial modes
Solar physicists routinely measure the horizontal flow field $\mathbf{u}_{h}(\textbf{{r}},t)$ at the solar surface, in the rotating frame $\Omega_{\rm ref}$ at intermediate spatial scales with a typical temporal cadence of one day [GizonInertial21]. These measurements are obtained using local correlation tracking (LCT), which infers large-scale flows from the motion of small-scale convective structures (granules) observed in intensity images. High-resolution imaging from instruments such as HMI on the SDO spacecraft is required to resolve these granules. The stream function can be computed from $\Psi=-\Delta_{h}^{-1}(\operatorname{curl_{\mathrm{h}}}(\rho_{0}\mathbf{u}_{h})),$ then Fourier transformed in time and longitude to obtain $\Psi^{m,\omega}(\theta)$, which provides data for the inverse problems. Due to the limited observation coverage (only half of the Sun is visible at any time), the azimuthal Fourier transform integrates over $0,\pi]$ instead of $0,2\pi]$, causing leakage of the signal between the harmonic degrees $m$; see [Basu2016] for a review. We will not concern ourselves with this aspect, but instead focus on the observational limitation: higher latitudes cannot easily be observed due to the line-of-sight projection. Thus, we assume to observe
$y^{\delta}(\theta):=\Psi^{m,\omega}(\theta)+\mathrm{noise}(\theta),\qquad\theta\in I:=(0,\pi)\qquad\text{or}\qquad\theta\in I^{\prime}:=(\epsilon,\pi-\epsilon),\epsilon\geq 0,$ (14)
where $\delta=\|\mathrm{noise}\|$ denotes the noise level. The case $\epsilon=0$ corresponds to full surface coverage, while $\epsilon>0$ takes into account the missing data at higher latitudes.
@@ 3 Inertial waves on the manifold $r{\mathbb{S}^{2}}$}
With the model established, we now address the forward problem by proving existence, uniqueness and stability of wave solutions, for both the full equation (11) and its separated counterpart (12). We first show the well-posedness of Eq. (11) associated to the differential operator
@@
with recalling that $\beta_{\Omega}(\theta)=\Omega(\theta)-\Omega_{\mathrm{ref}}$ and $\alpha_{\Omega}(\theta)=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\big(\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\big)$.
Consider the space $H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ of functions in $H^{2}({r{\mathbb{S}^{2}}})$ with zero mean defined in Appendix A. By partial integration with taking into account symmetry of the bi-Laplacian and Laplacian operators (see Appendix, Lemma 19), the corresponding sesquilinear form $\mathbf{A}:H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\times H_{\diamond}^{2}({r{\mathbb{S}^{2}}})\to\mathbb{C}$ defined as $\mathbf{A}[\Psi,\psi]:=\langle\mathbf{B}(\Psi),\psi\rangle_{H_{\diamond}^{-2},H_{\diamond}^{2}}$ is given by
$\begin{aligned} \mathbf{A}[\Psi,\psi]&=\gamma\left\langle\Delta_{\mathrm{h}}\Psi,\Delta_{\mathrm{h}}\psi\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\Psi,\operatorname{\mathbf{grad}_{\mathrm{h}}}\psi\right\rangle+(\mathbf{A}^{\alpha}+\mathbf{A}^{\beta})[\Psi,\psi],\\ &\mathbf{A}^{\alpha}[\Psi,\psi]:=-\left\langle\alpha_{\Omega}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle,\hskip 18.49988pt\mathbf{A}^{\beta}[\Psi,\psi]:=\left\langle\beta_{\Omega}\Delta_{\mathrm{h}}\Psi,\frac{\partial\psi}{\partial\phi}\right\rangle.\end{aligned}$ (16)
@@
$\qquad=\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle-i\omega\left\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi_{m},\operatorname{\mathbf{grad}_{\mathit{m}\!}}\psi_{m}\right\rangle+(\mathbf{A}^{\alpha}_{m}+\mathbf{A}^{\beta}_{m})[\Psi_{m},\psi_{m}],$
$\mathbf{A}^{\alpha}_{m}[\Psi_{m},\psi_{m}]:=im\left\langle\alpha_{\Omega}\Psi_{m},\psi_{m}\right\rangle,\hskip 18.49988pt\;\mathbf{A}^{\beta}_{m}[\Psi_{m},\psi_{m}]:=-im\left\langle\beta_{\Omega}\Delta_{m}\Psi_{m},\psi_{m}\right\rangle.$
with $\mathbf{A}^{\alpha}_{m},\mathbf{A}^{\beta}_{m}:V_{m}\times V_{m}\to\mathbb{C}$ as restrictions of $\mathbf{A}^{\alpha}$, $\mathbf{A}^{\beta}$ defined in (16). Here, $\Delta_{m}$, $\operatorname{\mathbf{grad}_{\mathit{m}\!}}$ are the decomposed operators as defined in (61), (62).
@@ 3.1 Boundedness
It is straightforward to verify that $\mathbf{A}^{\alpha}$ and $\mathbf{A}^{\alpha}_{m}$ are bounded if $\Omega\in W^{2,\infty}(0,\pi)$. However, to reduce the regularity assumption on the unknown rotation, we only constrain
@@
Under this reduced regularity assumption, establishing boundedness requires additional technical analysis. In the following, we will frequently employ continuous embeddings between function spaces $X,Y$ denoted the embedding constants $C_{X\to Y}$, i.e., $\|\cdot\|_{Y}\leq C_{X\to Y}\|\cdot\|_{X}$.
@@ 3.2 Fredholm property and well-posedness
Since the embeddings $H^{1}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ and $H^{3/2}_{\diamond}({r{\mathbb{S}^{2}}})\hookrightarrow H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ are compact, $\mathbf{B}$ is a Fredholm operator of index 0; see, e.g., [Renardy1992, Sec. 8.2.4].
@@
$\qquad\geq\frac{\gamma}{2}\|\Psi\|^{2}_{H_{\diamond}^{2}}+\left(|\omega|\!-\!\frac{4}{\gamma^{3}}(C_{H_{\diamond}^{1}\to L^{6}}C_{H_{\diamond}^{1/2}\to L^{3}})^{4}\big(\|\Omega\!-\!\Omega_{\mathrm{{ref}}}\|^{2}_{H^{1}}\!+\!9\|\Omega\|^{2}_{H^{1}}\big)^{2}\right)\|\Psi\|^{2}_{H_{\diamond}^{1}}.$
If $|\omega|$ satisfies (26), then $\mathbf{A}$ is coercive, so $\mathbf{B}$ is boundedly invertible by the Lax-Milgram theorem [Evans, Section 6.2, Theorem 1]. ∎
@@
The elements of $\operatorname{ker}\mathbf{B}_{\omega}$ for $\omega\in\Lambda$ are called inertial modes and have been studied intensively in [Damien22] and other publications.
@@
Therefore, we are in the framework of the analytic Fredholm theorem (see [Renardy1992, Thm. 7.92]). Note that $\mathbf{B}_{\omega}$ is boundedly invertible if and only of if $I+\mathbf{C}(\omega)$ is boundedly invertible and that the kernels of both operators coincide. The first alternative of this theory, that is $(I+\mathbf{C}(\omega))^{-1}$ does not exist for any $\omega\in\mathbb{C}$, is excluded by the second part of Theorem 3. Therefore, the second alternative holds true and provides the statements of this corollary. ∎
@@ 3.3 Separated equations
$\left\langle\mathbf{B}_{m}\Psi_{m},\psi_{m}\right\rangle=\mathbf{A}_{m}[\Psi_{m},\psi_{m}],\hskip 18.49988pt\Psi_{m},\psi_{m}\in V_{m}.$ (27)
@@
$\qquad\leq|m|\left|\Im\big\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega}\Psi)\big\rangle\right|=|m|\left|\Im\big\langle\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi,\operatorname{\mathbf{grad}_{\mathit{m}\!}}(\beta_{\Omega})\Psi\big\rangle\right|$
$\qquad\leq|m|\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\Psi\|_{L^{6}}\|\operatorname{\mathbf{grad}_{\mathit{m}\!}}\beta_{\Omega}\|_{L^{2}}\|\Psi\|_{L^{3}}\leq\frac{|m|}{r}C_{H^{2}_{\diamond}\to L^{3}_{\diamond}}C_{H^{1}_{\diamond}\to L^{6}_{\diamond}}\|\Omega^{\prime}\|_{L^{2}({r{\mathbb{S}^{2}}})}\|\Psi\|_{H^{2}_{\diamond}}^{2}.$